If a person can walk 6 miles an hour, how long will it take for him to travel between his house and his job which is 10 miles away?
Answer by stanbon(75887)

If a person can walk 6 miles an hour, how long will it take for him to travel between his house and his job which is 10 miles away?
-----
time = distance/rate
---
t = 10/6 = 5/3 hr = 1 2/3 hrs = 1hr 40 min
==================================
Cheers,
Stan H.
=================